# mWhisper-Flamingo for Multilingual Audio-Visual Noise-Robust Speech Recognition

### v2

@@ Abstract
Audio-Visual Speech Recognition (AVSR) combines lip-based video with audio and can improve performance in noise, but most methods are trained only on English data. One limitation is the lack of large-scale multilingual video data, which makes it hard hard to train models from scratch. In this work, we propose mWhisper-Flamingo for multilingual AVSR which combines the strengths of a pre-trained audio model (Whisper) and video model (AV-HuBERT). To enable better multi-modal integration and improve the noisy multilingual performance, we introduce decoder modality dropout where the model is trained both on paired audio-visual inputs and separate audio/visual inputs. mWhisper-Flamingo achieves state-of-the-art WER on MuAViC, an AVSR dataset of 9 languages. Audio-visual mWhisper-Flamingo consistently outperforms audio-only Whisper on all languages in noisy conditions.
@@ III-B Clean Results
Next, we establish baselines using Whisper models. Compared to XLAVS-R 2B, Whisper small zero-shot achieves better performance on the lower resource languages (37.5% vs 41.9%), while Whisper medium zero-shot achieves better overall average non-En WER (22.9% vs 26.4%). This shows Whisper’s strong multilingual capabilities and motivates its selection as the foundation of our proposed models.
@@ III-C Noisy Results
Table II shows the performance on MuAViC with babble noise at 0-SNR. The babble noise is from Whisper-Flamingo [17] and was constructed from 30 LRS3 speakers. We compare Whisper zero-shot with Whisper fine-tuned and mWhisper-Flamingo. The noisy results show a significant performance degredation compared to the clean results (Table I). For example, the average non-En WER for Whisper small zero-shot increases from 27.0% in the clean setting to 71.0% under 0-SNR babble noise. However, fine-tuning significantly improves WER. For the small model, fine-tuning improves the average non-En WER from 71.0% (zero-shot) to 55.3%. A similar trend is observed for the medium model, with WER decreasing from 60.7% to 48.0%.
@@
mWhisper-Flamingo medium achieves the best multilingual WER (43.7%), even outperforming Whisper large zero-shot (65.0%) despite having fewer parameters (1.39B vs 1.5B). This shows that the performance gains are not only due to more parameters but also due to the visual modality. As additional evidence, mWhisper-Flamingo small (651M parameters) outperforms audio-only fine-tuned Whisper medium (769M parameters) on the higher-resource languages (37.4% vs. 38.1%). This highlights that smaller audio-visual models can achieve better performance than larger audio-only models.

### v3

@@ Abstract
Audio-Visual Speech Recognition (AVSR) combines lip-based video with audio and can improve performance in noise, but most methods are trained only on English data. One limitation is the lack of large-scale multilingual video data, which makes it hard to train models from scratch. In this work, we propose mWhisper-Flamingo for multilingual AVSR which combines the strengths of a pre-trained audio model (Whisper) and video model (AV-HuBERT). To enable better multi-modal integration and improve the noisy multilingual performance, we introduce decoder modality dropout where the model is trained both on paired audio-visual inputs and separate audio/visual inputs. mWhisper-Flamingo achieves state-of-the-art WER on MuAViC, an AVSR dataset of 9 languages. Audio-visual mWhisper-Flamingo consistently outperforms audio-only Whisper on all languages in noisy conditions.
@@ III-B Clean Results
Next, we establish baselines using Whisper. Compared to XLAVS-R 2B, Whisper small zero-shot (without fine-tuning) achieves better performance on the lower resource languages (37.5% vs 41.9%), while Whisper medium zero-shot achieves better overall average non-En WER (22.9% vs 26.4%). This shows Whisper’s strong multilingual capabilities and motivates its selection as the foundation of our proposed models.
@@ III-C Noisy Results
Table II shows the performance on MuAViC with babble noise at 0-SNR. The babble noise is from Whisper-Flamingo [17] and was constructed from 30 LRS3 speakers. We compare Whisper zero-shot with Whisper fine-tuned and mWhisper-Flamingo. The noisy results show a significant performance degradation compared to the clean results (Table I). For example, the average non-En WER for Whisper small zero-shot increases from 27.0% in the clean setting to 71.0% under 0-SNR babble noise. However, fine-tuning significantly improves WER. For the small model, fine-tuning improves the average non-En WER from 71.0% (zero-shot) to 55.3%. A similar trend is observed for the medium model, with WER decreasing from 60.7% to 48.0%.
@@
mWhisper-Flamingo medium achieves the best multilingual WER (43.7%), even outperforming Whisper large zero-shot (56.0%) despite having fewer parameters (1.39B vs 1.5B). This shows that the performance gains are not only due to more parameters but also due to the visual modality. As additional evidence, mWhisper-Flamingo small (651M parameters) outperforms audio-only fine-tuned Whisper medium (769M parameters) on the higher-resource languages (37.4% vs. 38.1%). This highlights that smaller audio-visual models can achieve better performance than larger audio-only models.